# SDC-Net: A Domain Adaptation Framework with Semantic-Dynamic Consistency for Cross-Subject EEG Emotion Recognition

### v1

@@ III-B Dynamic Distribution Alignment in RKHS
To address the distributional discrepancy between source domain $D_{s}$ and target domain $D_{t}$ in EEG emotion recognition, we propose a unified framework named Dynamic Distribution Alignment in RKHS. This method performs joint alignment of MPD and CPD in a shared RKHS. The key idea is to unify marginal and conditional alignment into a single kernel mean embedding (KME)-based framework.
III-B1 Unified Alignment Framework
Given a kernel function $k(x,x^{\prime})=\langle\psi(x),\psi(x^{\prime})\rangle$, the KME of a distribution $P(x)$ is:
$\mu_{P}=\mathbb{E}_{x\sim P}[\psi(x)]$ (6)
For a conditional distribution $P(x|y)$, the embedding is:
$\mu_{X|y}=\mathbb{E}_{X|y}[\psi(X)]=M_{X|Y}\varphi(y)$ (7)
where $M_{X|Y}$ is the conditional embedding operator, and $\varphi(y)$ denotes the label kernel feature. This leads to the following unified alignment objective:
$\mathcal{L}_{\text{align}}=\mathbb{E}_{y\sim Y}\left\|\mu_{P(x|y)}-\mu_{Q(x|y)}\right\|_{\mathcal{H}}^{2}$ (8)
This unified objective naturally encompasses both marginal and conditional alignment scenarios. When label information is not considered and only the marginal distribution $P(x)$ is aligned, we obtain the standard MMD for aligning global feature distributions:
$\mathcal{L}_{\text{MMD}}=\left\|\mu_{P(x)}-\mu_{Q(x)}\right\|_{\mathcal{H}}^{2}$ (9)
When class-conditional structures are taken into account, the objective becomes the CMMD, which aligns distributions conditioned on class labels.
$\mathcal{L}_{\text{CMMD}}=\sum_{c=1}^{C}\left\|\mu_{P(x|y=c)}-\mu_{Q(x|y=c)}\right\|_{\mathcal{H}}^{2}$ (10)
III-B2 Alignment of Marginal Distribution Alignment In RKHS
Given source domain samples $D_{s}=\{x_{i}^{s}\}_{i=1}^{n}$ and target domain samples $D_{t}=\{x_{j}^{t}\}_{j=1}^{m}$, the empirical MMD is computed as:
$\mathcal{L}_{\text{MMD}}$ $=\frac{1}{n^{2}}\sum_{i,j=1}^{n}k(x_{i}^{s},x_{j}^{s})+\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(x_{i}^{t},x_{j}^{t})$ (11)
$\quad-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}k(x_{i}^{s},x_{j}^{t})$
We adopt a multi-kernel Gaussian kernel to enhance robustness:
$\begin{split}k(u,v)&=\sum_{l=1}^{L}\exp\left(-\frac{\|u-v\|^{2}}{\sigma_{l}}\right)\\ \sigma_{l}&=\sigma_{0}\cdot\kappa^{l-L/2}\end{split}$ (12)
where $\kappa$ is the kernel multiplier, $L$ is the number of Gaussian kernels, and $\sigma_{0}$ is the base bandwidth computed as the average squared pairwise distance between all samples. This formulation enables capturing similarities at multiple scales and improves the robustness of domain discrepancy estimation.
III-B3 Alignment of Conditional Probability Distributions In RKHS
Most prior work on domain adaptation in EEG-based emotion recognition focuses on aligning marginal distributions between the source and target domains. However, aligning the CPDs, i.e., $P(x|y)$, is equally critical to ensure semantic consistency across domains. To address this, we adopt the CMMD framework, which measures the discrepancy between class-conditional distributions in a RKHS.
The CMMD loss is defined as:
$\mathcal{L}_{\text{CMMD}}=\sum_{c=1}^{C}\left\|\mu_{P(x|y=c)}-\mu_{Q(x|y=c)}\right\|_{\mathcal{H}}^{2}$ (13)
where $\mu_{P(x|y=c)}$ and $\mu_{Q(x|y=c)}$ denote the KME of the conditional distributions from the source and target domains, respectively.
Since the target domain lacks ground-truth labels, we use pseudo-labels predicted by the classifier: $\hat{y}_{t}^{u}=f(x_{t}^{u})$. Rather than estimating posteriors $P(y|x)$ and $Q(y|x)$ directly, we focus on sufficient statistics of the class-conditional distributions $P(x|y)$ and $Q(x|y)$.
In RKHS, the conditional mean embedding is expressed as Eq.7. The operator $M_{X|Y}$ can be constructed from the covariance operators:
$M_{X|Y}=\mathcal{S}_{XY}\mathcal{S}_{YY}^{-1}$ (14)
with the cross-covariance operator defined as:
$\mathcal{S}_{XY}=\mathbb{E}_{XY}[\psi(X)\otimes\varphi(Y)]-\mathbb{E}_{X}[\psi(X)]\otimes\mathbb{E}_{Y}[\varphi(Y)]$ (15)
where $\otimes$ denotes the tensor product.
In practical settings, the conditional embedding operator can be estimated from $m$ samples. Define the input and label feature matrices:
$\Theta$ $=[\psi(x_{1}),\dots,\psi(x_{m})]\in\mathbb{R}^{d\times m},$ (16)
$\xi$ $=[\varphi(y_{1}),\dots,\varphi(y_{m})]\in\mathbb{R}^{k\times m}$ (17)
We compute the centered empirical cross-covariance:
$\hat{S}_{XY}=\frac{1}{m}\Theta H\xi^{\top},\quad H=I-\frac{1}{m}\mathbf{1}\mathbf{1}^{\top}$ (18)
The estimated conditional embedding operator is then:
$\hat{M}_{X|Y}=\Theta(\xi^{\top}\xi+\lambda I)^{-1}\xi^{\top}$ (19)
where $\lambda$ is a small regularization coefficient to ensure numerical stability.
Let the feature and label matrices from source and target domains be:
  • Source domain: $\Theta_{s}$, $\xi_{s}$, kernel matrices $K_{s}=\Theta_{s}^{\top}\Theta_{s}$, $\Upsilon_{s}=\xi_{s}^{\top}\xi_{s}$
  • Target domain: $\Theta_{t}$, $\xi_{t}$, kernel matrices $K_{t}=\Theta_{t}^{\top}\Theta_{t}$, $\Upsilon_{t}=\xi_{t}^{\top}\xi_{t}$
  • Cross-domain: $K_{st}=\Theta_{s}^{\top}\Theta_{t}$, $\Upsilon_{ts}=\xi_{t}^{\top}\xi_{s}$
Define regularized matrices:
$\tilde{\Upsilon}_{s}=\Upsilon_{s}+\lambda I,\quad\tilde{\Upsilon}_{t}=\Upsilon_{t}+\lambda I$
Then the CMMD loss in matrix form becomes:
^CMMD=subscript^$\widehat{\mathcal{L}}_{\text{CMMD}}=$ $\operatorname{Tr}\left(\Upsilon_{s}\tilde{\Upsilon}_{s}^{-1}K_{s}\tilde{\Upsilon}_{s}^{-1}\right)+\operatorname{Tr}\left(\Upsilon_{t}\tilde{\Upsilon}_{t}^{-1}K_{t}\tilde{\Upsilon}_{t}^{-1}\right)$ (20)
$-2\cdot\operatorname{Tr}\left(\Upsilon_{ts}\tilde{\Upsilon}_{s}^{-1}K_{st}\tilde{\Upsilon}_{t}^{-1}\right)$
@@ III-B3 Alignment of Conditional Probability Distributions In RKHS
To mitigate the effect of noisy pseudo-labels from the target domain, we adopt a dynamic confidence-based selection mechanism. Let $\hat{\mathbf{y}}_{i}^{t}$ denote the predicted label distribution for target sample $\mathbf{x}_{i}^{t}$. If the confidence $\max(\hat{\mathbf{y}}_{i}^{t})\geq\tau$, the pseudo-label is accepted; otherwise, the sample is excluded from CMMD computation. Formally:
@@
The threshold $\tau$ increases over training to incorporate more reliable samples. The final CMMD loss (Eq. 20) is computed only on high-confidence samples from $\hat{D}_{t}$ to ensure robust conditional distribution alignment.
Figure 3: The flowchart of the proposed SDC-Net framework.
III-C Dual-Domain Similarity Consistency Learning Strategy
To promote discriminative and transferable feature representations for EEG signals, we propose a DSCL strategy. This module encourages the learned feature space to maintain pairwise semantic consistency—ensuring that samples from the same class are closer, and those from different classes are further apart. It does so by applying pairwise similarity learning in both the source and target domains, as illustrated in Fig. 3.
III-C1 Supervised Similarity Consistency on Source Domain
@@
III-C2 Self-Inferred Similarity Consistency on Target Domain
In the target domain $D_{t}$, labels are unavailable. To still enable pairwise semantic learning, we utilize pseudo-labels $\hat{y}_{t}$ generated from the model itself using the confidence-based filtering strategy (see Eq. LABEL:eq:pseudo_label).
@@ III-C2 Self-Inferred Similarity Consistency on Target Domain
Here, $\tau_{pu}$ and $\tau_{pl}$ represent dynamic upper and lower thresholds used to determine confident positive or negative pairs, respectively. Feature similarities within the ambiguous region $[\tau_{pl},\tau_{pu}]$ are excluded to avoid noisy supervision. These thresholds are linearly adjusted during training to gradually include more pairs as the model becomes more confident (visualized in Fig. 3).
@@
III-D Loss Function and Training Procedure of SDC-Net
@@
where $B_{L}$ is the batch size, $C$ denotes the number of emotion classes, $y_{s}^{l}$ is the one-hot ground-truth label, and $p_{s}^{l}$ is the predicted probability distribution.
@@
This progressive adjustment ensures that SDC-Net shifts from global alignment to finer semantic refinement as training evolves, ultimately improving generalization performance in fully unsupervised cross-subject EEG emotion recognition tasks. The overall learning process is detailed in Algorithm 1.
Algorithm 1 SDC-Net: Domain-Adaptive Representation Learning
0:  Source domain dataset $D_{s}=\{(x_{s}^{l},y_{s}^{l})\}_{i=1}^{N}$, Target domain dataset $D_{t}=\{x_{t}^{u}\}_{i=1}^{N_{u}}$, Percentile parameter $\alpha$, Total epochs $E$
0:  Predicted labels for target domain $D_{t}$
1:  for epoch $e=1$ to $E$ do
2:     Sample interpolation weight $\omega\sim\text{Beta}(\alpha,\alpha)$. Augment trial data via convex combinations per Eqs. (3), (4).
3:     Extract mini-batches $D_{\text{batch}}^{s}\subseteq D_{s}$ and $D_{\text{batch}}^{t}\subseteq D_{t}$ through random sampling.
4:     Derive 310-dimensional EEG feature representations using short-time Fourier transform and differential entropy.
5:     Compute MMD loss to align marginal distributions, as per Eq. (11).
6:     Infer pseudo-labels $\hat{y}_{t}^{u}=f(x_{t}^{u})$ and probability distributions $p_{i}^{u}$ for $x_{t}^{u}\in D_{\text{batch}}^{t}$, addressing the absence of ground-truth labels in $D_{t}$.
7:     Compute dynamic confidence threshold. Select high-confidence samples as per Eq. 21.
8:     Evaluate CMMD loss using $\hat{D}_{t}$ per Eq. (20), and compute source domain cross-entropy loss per Eq. (28).
9:     Calculate pairwise alignment losses for $D_{s}$ and $D_{t}$ using cosine similarity and thresholded pseudo-labels, as defined in Eqs. (24), (26).
10:     Optimize network parameters by minimizing the composite loss $\mathcal{L}$ with dynamically adjusted weights $\alpha$, $\beta$, $\lambda$ per Eqs. (27), (29), (30), via gradient descent.
11:  end for
12:  return  Predicted labels for $D_{t}$.
IV Experiments
IV-A Emotion datasets
To demonstrate the effectiveness of our method, SDC-Net for cross-subject EEG-based emotion recognition, we conduct experiments on four public benchmark datasets: SEED, SEED-IV, and FACED.
In the SEED dataset, fifteen Chinese subjects participated in the experiments, with each subject completing three sessions, each consisting of a total of 15 trials. Throughout the experiments, participants were exposed to Chinese film clips specifically designed to elicit a spectrum of emotions, including positive, neutral, and negative states[30].
In the SEED IV dataset, a total of 15 subjects took part in the experiments. For each participant, three sessions were conducted on separate days, with each session comprising 24 trials. During each trial, the participant viewed one of several film clips designed to induce emotions such as happiness, sadness, neutrality, and fear. The EEG signals in both the SEED and SEED IV datasets were recorded using a 62-channel ESI NeuroScan System[31].
In the FACED dataset, a total of 123 Chinese participants took part in the experiments, each completing 28 trials. Four trials were conducted to elicit neutral emotions, while the remaining trials targeted various other emotions. The movie clips lasted between 34 and 129 seconds and were specifically selected to induce a range of emotions, including anger, fear, disgust, sadness, amusement, inspiration, joy, tenderness, and neutrality. During the emotion induction process, EEG signals were recorded using the NeuSen.W32 system (Neuracle, China) with 32 channels. Two types of classification tasks were performed on the FACED dataset: one for fine-grained discrete emotion classification (nine categories) and another for binary classification to distinguish between positive emotions (amusement, inspiration, joy, tenderness) and negative emotions (anger, fear, disgust, sadness)[32].
IV-B Implementation details
In the model presented in this paper, the feature extractor is composed of a pair of fully connected layers. The initial layer, denoted as ”fc1,” processes input data spanning 310 dimensions and reduces it to 64 dimensions using a ReLU activation function. Subsequently, the data flows through the second layer, labeled ”fc2,” further reducing the dimensionality to 64, accompanied by an additional ReLU activation. To optimize the model’s performance and mitigate overfitting, dropout layers with a dropout rate of 0.25 are introduced after each fully connected layer. For the weight parameters of $\alpha$, $\tau_{h}$ and $\tau_{l}$ are 1 and 0, respectively, while $\tau_{m}=\frac{2}{1+e^{-10\cdot\frac{\text{e}}{\text{200}}}}-1$, which is a logistic sigmoid-based function that transitions smoothly and tends to zero as the epoch progresses. For the Eq. (29) , $\rho_{0}$ and $\rho_{1}$ are 0.1 and 0.15, respectively. The parameters $\tau_{pu}$ and $\tau_{pu}$ in pairwise learning are assigned values of 0.9 and 0.5, respectively.
For our experimental setup, we utilize all labeled source samples to represent the $D_{s}$, while the unlabeled target samples constitute the $D_{t}$. We evaluate and compare the average classification accuracy as a performance measure. Our training configuration encompasses several essential parameters and procedural steps. The batch size for training is set to 32, and the training process spans 200 epochs. We employ two distinct learning rates, specifically 0.001 and 0.01. The SGD momentum parameter is set to 0.9. Moreover, the option to enable or disable CUDA training is provided to the user. Additionally, we initialize a random seed value of 3 to ensure result reproducibility, and the L2 weight decay is precisely set at $5\times 10^{-4}$. During training, we employ the SGD optimization method, with learning rates subject to dynamic adjustments at the beginning of each epoch. These adaptive changes in learning rates follow a specific mathematical formula. This iterative training process continues for the designated number of epochs, allowing the model to learn and adapt over time.
IV-C Experiment Setting
In this study, the SEED and SEED_IV datasets were utilized as the foundational data sources for model training and evaluating the performance of cross-subject emotion detection. To comprehensively assess the effectiveness of the cross-subject emotion detection model, two distinct cross-validation methodologies were implemented. These methodologies aimed to provide a robust evaluation of the model’s performance across different subjects.
Cross-Subject Single-Session Leave-One-Subject-Out Cross-Validation Results. This approach employs a widely embraced technique for cross-subject emotion detection using EEG signals. Specifically, a single-session leave-one-subject-out cross-validation strategy is employed. In our experiments, the EEG signals of each individual subject were designated as the $D_{t}$, while the data from all other subjects constituted the $D_{s}$. This process was systematically repeated, with each subject serving as the $D_{t}$ in turn. The cross-subject average detection performance was then reported, providing a comprehensive evaluation of the approach’s effectiveness.
Cross-Subject Cross-Session Leave-One-Subject-Out Cross-Validation Results. To comprehensively evaluate the model’s performance across different sessions for each subject, we employed a cross-subject, cross-session leave-one-subject-out cross-validation method. In each experimental iteration, the complete session data of a specific subject was used as the $D_{t}$, while the sessions from all other subjects constituted the $D_{s}$. This iterative procedure ensured that each subject served as the $D_{t}$ at least once, enabling the calculation of the average detection performance across subjects.
@@ IV-C Experiment Setting
Method Pcc(%) Method Pcc(%)
RGNN[33] 85.30$\pm$06.72 BiHDM[34] 85.40$\pm$07.53
JDA-Net[9] 88.28$\pm$11.44 DA-CapsNet[35] 84.63$\pm$09.09
MS-MDA[7] 89.63$\pm$06.97 WGAN-GP[36] 87.10$\pm$07.10
DGGN[37] 83.84$\pm$10.26 EPNNE[12] 89.10$\pm$03.60
DC-ASTGCN[38] 80.65$\pm$08.46 MS-FRAN[8] 85.61$\pm$06.55
SDDA[11] 91.08$\pm$07.70 CU-GCN[39] 87.10$\pm$05.44
DAPLP [10] 89.44$\pm$04.89 DS-AGC [40] 86.38$\pm$07.25
PR-PL[13] 85.88$\pm$09.36 PLMSDANet[41] 84.21$\pm$12.34
CLISA[15] 86.40$\pm$06.40 CL-CS[17] 88.30$\pm$08.90
ST-SCGNN[42] 85.90$\pm$04.90 SDC-Net 91.85$\pm$05.98
@@
TABLE III: The performance of representation methods on SEED_IV datasets using Cross-Subject Single-Session Leave-One-Subject-Out Cross-Validation
Method Pcc(%) Method Pcc(%)
DGCNN[43] 68.73$\pm$08.34 MS-STM[44] 61.41$\pm$09.72
MS-ADRT[45] 68.98$\pm$06.80 MS-MDA[7] 59.34$\pm$05.48
WGAN-GP[36] 60.60$\pm$15.76 JDA-Net[9] 70.83$\pm$10.25
ST-SCGCN[42] 76.37$\pm$05.77 CU-GCN[39] 74.50$\pm$07.88
DAPLP [10] 74.57$\pm$06.18 DS-AGC [40] 66.00$\pm$07.93
MSFR-GCN[46] 73.43$\pm$07.32 SDC-Net 74.88$\pm$10.47
TABLE IV: Performance Comparison of Methods on FACED-2 and FACED-9 Datasets
Method FACED-2 (Pcc%) FACED-9 (Pcc%)
DE+SVM[47] 60.50$\pm$16.00 34.90$\pm$10.70
DE+MLP[48] 70.20$\pm$11.20 35.10$\pm$10.30
DAN[49] 53.40$\pm$07.10 37.20$\pm$09.10
DANN[50] 50.50$\pm$07.10 54.10$\pm$08.30
CLISA[15] 67.80$\pm$04.10 43.20$\pm$05.90
GCPL[51] / 36.90$\pm$03.30
CL-CS[17] 72.50$\pm$15.30 43.40$\pm$13.70
SDC-Net (Ours) 75.20$\pm$08.46 42.40$\pm$06.55
@@
TABLE V: The performance of shared methods on SEED and SEED_IV datasets using Cross-Subject Cross-Session Leave-One-Subject-Out Cross-Validation
Method SEED SEED_IV
RF[52] 69.60$\pm$07.64 50.98$\pm$09.20
KNN[53] 60.66$\pm$07.93 40.83$\pm$07.28
SVM[54] 68.15$\pm$07.38 51.78$\pm$12.85
TCA[55] 64.02$\pm$07.96 56.56$\pm$13.77
CORAL[56] 68.15$\pm$07.83 49.44$\pm$09.09
SA[57] 61.41$\pm$09.75 64.44$\pm$09.46
GFK[58] 66.02$\pm$07.59 45.89$\pm$08.27
DANN[59] 81.08$\pm$05.88 54.63$\pm$08.03
SDC-Net 82.22$\pm$04.68 68.84$\pm$08.05
Figure 4: The result of nine categories for Faced datasets.
(a) DA-CapsNet
(b) PLMSDANet
(c) LGDAAN-Net
(d) SDC-Net
Figure 5: Confusion matrices comparison: (a) DA-CapsNet[35], (b) PLMSDANet[41], (c) LGDAAN-Net[60], and (d) proposed SDC-Net. All subfigures share the same color scale for fair comparison.
V Result and Discussion
V-A Experimental Results
V-A1 Cross-subject single-session leave-one-out-subject-out cross-validation Results
In Tables II and III, we conducted a comprehensive evaluation of various representations on the SEED and SEED_IV datasets using the leave-one-subject-out cross-validation method with a cross-subject single-session protocol. Our method demonstrated significant performance advantages on the SEED dataset, achieving an accuracy of 91.85%$\pm$ 05.98%. Similarly, on the SEED_IV dataset, our method exhibited competitive accuracy of 74.88%$\pm$10.47%. These results strongly validate the substantial performance improvements achieved by our method on both datasets, particularly the remarkable accuracy of 91.85%$\pm$ 05.98% on the SEED dataset, surpassing the industry average and demonstrating notable potential in the field of emotion recognition tasks. These findings provide compelling evidence supporting the effective application of our method in real-world scenarios.
As shown in Table IV, the SDC-Net model was evaluated on the FACED dataset for both binary classification of positive and negative emotions and nine-class emotion classification, achieving accuracies of 75.2%$\pm$8.46and 42.4%$\pm$6.55%, respectively. Compared to the baseline model DE±$\pm$±SVM, it achieved improvements of and for the binary and nine-class tasks, respectively, highlighting the superior performance of our model, particularly in handling more fine-grained emotion classification tasks.
V-A2 Cross-subject cross-session leave-one-out-subject-out cross-validation results
Another crucial consideration for emotion brain-computer interfaces is the substantial variability observed among different subjects across various sessions. The evaluation approach of cross-subject and cross-session represents a significant challenge for the effectiveness of models in EEG-based emotion recognition tasks. To further validate this detection approach, which aligns more closely with real-world application scenarios, we conducted experiments and obtained outstanding three-class classification performance on the SEED dataset, achieving an accuracy of 82.22%$\pm$04.68(see Table V). Additionally, on the SEED-IV dataset, our model achieved a four-class accuracy of 68.84%$\pm$08.05(see Table V). Compared to existing research, the proposed SDC-Net method demonstrated industry-leading performance with a smaller standard deviation. These results indicate that the proposed SDC-Net method exhibits excellent stability and generalization capabilities in handling subject and session differences.
V-B Confusion matrices

### v2

@@ II-B Dynamic Distribution Alignment in RKHS
To address the distributional discrepancy between source domain $D_{s}$ and target domain $D_{t}$ in EEG emotion recognition, we propose a unified framework named Dynamic Distribution Alignment in RKHS. This method performs joint alignment of MPD and CPD in a shared RKHS. The key idea is to unify marginal and conditional alignment into a single kernel mean embedding (KME)-based framework. The detailed derivations of Dynamic Distribution Alignment in RKHS are provided in the Supplementary Information (Section S1).
II-B1 Unified Alignment Framework
We adopt a unified kernel mean embedding (KME) framework to align both marginal and conditional distributions in the RKHS. Formally, the alignment loss is defined as
$\mathcal{L}_{\text{align}}=\mathbb{E}_{y\sim Y}\|\mu_{P(x|y)}-\mu_{Q(x|y)}\|_{\mathcal{H}}^{2},$ (6)
where $\mu_{P(x|y)}$ and $\mu_{Q(x|y)}$ denote the conditional embeddings of the source and target domains. Detailed derivations and special cases (MMD, CMMD) are provided in the Supplementary Information (Section S1.1).
II-B2 Alignment of Marginal Distributions in RKHS
To reduce global distribution shift, we employ the Maximum Mean Discrepancy (MMD) between source and target domains:
$\mathcal{L}_{\text{MMD}}=\big\|\tfrac{1}{n}\!\sum_{i=1}^{n}\varPhi(x_{i}^{s})-\tfrac{1}{m}\!\sum_{j=1}^{m}\varPhi(x_{j}^{t})\big\|_{H}^{2},$ (7)
where $\varPhi(\cdot)$ denotes the feature mapping into RKHS. The empirical estimation and multi-kernel extension of MMD are derived in the Supplementary Information (Section S1.2).
II-B3 Alignment of Conditional Probability Distributions in RKHS
@@
To ensure robust conditional distribution alignment, we first introduce a dynamic confidence-based selection mechanism for pseudo-labels in the target domain. Specifically, for each target sample $\mathbf{x}_{i}^{t}$, let $\hat{\mathbf{y}}_{i}^{t}$ denote its predicted label distribution. If the confidence $\max(\hat{\mathbf{y}}_{i}^{t})\geq\tau$, the pseudo-label is accepted; otherwise, the sample is excluded from CMMD computation. Formally:
@@
The threshold $\tau$ is gradually increased during training, allowing more reliable samples to be incorporated over time. Consequently, the final CMMD loss is computed only on high-confidence samples from $\hat{D}_{t}$.
Beyond marginal alignment, we further enforce semantic consistency by aligning class-conditional distributions across domains. The CMMD loss is defined as
$\mathcal{L}_{\text{CMMD}}=\sum_{c=1}^{C}\|\mu_{P(x|y=c)}-\mu_{Q(x|y=c)}\|_{\mathcal{H}}^{2},$ (9)
where $\mu_{P(x|y=c)}$ and $\mu_{Q(x|y=c)}$ denote the mean embeddings of source and target samples for class $c$ in RKHS. Detailed derivations of CMMD are provided in the Supplementary Information (Section S1.3).
Figure 2: An illustration of the proposed DSCL strategy. The left part shows supervised similarity consistency learning on the source domain using ground-truth labels. The right part depicts self-inferred similarity consistency learning on the target domain based on pseudo-labels and dynamic similarity thresholds. Pairwise similarity is enforced through BCE loss in both domains to promote intra-class compactness and inter-class separation.
II-C Dual-Domain Similarity Consistency Learning Strategy
To promote discriminative and transferable feature representations for EEG signals, we propose a DSCL strategy. This module encourages the learned feature space to maintain pairwise semantic consistency—ensuring that samples from the same class are closer, and those from different classes are further apart. In order to achieve it, pairwise similarity learning is applied in both the source and target domains, as illustrated in Fig.2.
II-C1 Supervised Similarity Consistency on Source Domain
@@
II-C2 Self-Inferred Similarity Consistency on Target Domain
In the target domain $D_{t}$, labels are unavailable. To still enable pairwise semantic learning, we utilize pseudo-labels $\hat{y}_{t}$ generated from the model itself using the confidence-based filtering strategy.
@@ II-C2 Self-Inferred Similarity Consistency on Target Domain
Here, $\tau_{pu}$ and $\tau_{pl}$ represent dynamic upper and lower thresholds used to determine confident positive or negative pairs, respectively. Feature similarities within the ambiguous region $\tau_{pl},\tau_{pu}]$ are excluded to avoid noisy supervision. These thresholds are linearly adjusted during training to gradually include more pairs as the model becomes more confident (visualized in Fig.2).
@@
II-D Loss Function and Training Procedure of SDC-Net
@@
where $B_{L}$ is the batch size, $C$ denotes the number of emotion classes, $y_{s}^{l}$ presents the one-hot ground-truth label, and $p_{s}^{l}$ means the predicted probability distribution.
@@
This progressive adjustment ensures that SDC-Net shifts from global alignment to finer semantic refinement as training evolves, ultimately improving generalization performance in fully unsupervised cross-subject EEG emotion recognition tasks. The overall learning process is provided in Algorithm S1 of the Supplementary Information.
III Experiments
III-A Emotion datasets
To demonstrate the effectiveness of SDC-Net for cross-subject EEG-based emotion recognition, we conduct experiments on three public benchmark datasets: SEED [25], SEED-IV [26], and FACED [27]. SEED and SEED-IV contain EEG signals from 15 subjects recorded with a 62-channel NeuroScan system, induced by film clips designed to evoke three (positive, neutral, negative) and four (happiness, sadness, neutral, fear) emotions, respectively. FACED involves 123 participants with 32-channel recordings and supports both nine-class discrete emotion classification and a binary positive/negative task.
III-B Experiment Setting and Implementation Details
We evaluated SDC-Net on the SEED and SEED-IV datasets using two widely adopted cross-validation protocols: (1) Cross-subject single-session leave-one-subject-out, where one subject was used as the target domain $D_{t}$ and the others as the source domain $D_{s}$; (2) Cross-subject cross-session leave-one-subject-out, where one subject’s entire session was treated as $D_{t}$ and the remaining sessions as $D_{s}$. These protocols provide a rigorous assessment of generalization across both subjects and sessions. Further procedural details are provided in the Supplementary Information (Section S2.1).
The feature extractor consists of two fully connected layers (310 $\to$ 64 $\to$ 64) with ReLU activations and dropout (0.25). Training was performed for 200 epochs with SGD (momentum = 0.9), batch size = 32, and initial learning rates in $\{0.001,0.01\}$. Additional hyperparameter specifications (e.g., weight decay, random seed, threshold schedules, and dynamic coefficients) are reported in the Supplementary Information (Section S2.2-S2.4).
IV Result
IV-A Experimental Results
IV-A1 Cross-subject single-session leave-one-out-subject-out cross-validation Results
[…]
IV-A2 Cross-subject cross-session leave-one-out-subject-out cross-validation results
[…]
@@ IV-A2 Cross-subject cross-session leave-one-out-subject-out cross-validation results
Method Acc(%) Method Acc(%)
[…]
@@
TABLE III: The performance of representation methods on SEED-IV datasets using Cross-Subject Single-Session Leave-One-Subject-Out Cross-Validation
Method Acc(%) Method Acc(%)
[…]
@@
TABLE IV: The performance of shared methods on SEED and SEED-IV datasets using Cross-Subject Cross-Session Leave-One-Subject-Out Cross-Validation
Method SEED (Acc%) SEED-IV (Acc%)
[…]
IV-B Confusion matrices